OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 · x + 4 · 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10 @iGreen @Hotchellerae21 @monsterfox1997 @RyanTLopez

OpenStudy (anonymous):

First distribute 4 into the parenthesis: 4(3 + 7) = 4 * x + 4 * 7 12 + 28 = 4 * x + 4 * 7 Add: 40 = 4 * x + 4 * 7 Multiply: 40 = 4x + 28 Now subtract 40 - 28, what do you get?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yep, that gives us: 12 = 4x Now can you divide 12 / 4? @Wolf_Girly2002

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yep, so that's your answer!

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Wolf_Girly2002